OpenStudy (lexy148):

can someone help me solve this by using elimination or substitution: -4x+4y=4 5x-y=-13

OpenStudy (solomonzelman):

you can divide the first equation by 4, and then add the equations.

OpenStudy (lexy148):

y=2

OpenStudy (lexy148):

@solomonzelman

OpenStudy (solomonzelman):

i don't think so

OpenStudy (lexy148):

i moved 4 to one side then divided 8 by 4,which gave me 2

OpenStudy (solomonzelman):

y=2 is not correct

OpenStudy (solomonzelman):

-4x+4y=4 when you divide this equation by 4, you get - x + y = 1

OpenStudy (solomonzelman):

-x + y = 1 5x - y = -13 ------------- add the equations

OpenStudy (lexy148):

x=-3?

OpenStudy (retireed):

This is not meant as a criticism :) but you and the other students need to learn to check your own answer and not asking a mentor to verify it. How how you check your answer? First calculate y from the equation -x + y = 1 where you "suspect" x = -3 -(-3) + y = 1 3 + y = 1 y = 1 - 3 y = -2 Then plug x = -3 and y = -2 into the OTHER equation and see if it checks (is correct) 5x - y = -13 5(-3) - (-2) = -13 Is this correct? Check and see, but don't ask me.